### v3

# On Lan-Sheng-Zuo conjecture

Abstract.
In this paper we study the Lan-Sheng-Zuo conjecture proposed in [LSZ13]. We prove that the conjecture holds for smooth projective curves with genus $g\leq 1$, and construct explicit counter-examples of arbitrary big rank (the first example is $p=2,r=3$) Higgs bundles over any smooth projective curve with genus $g\geq 2$.
1. Introduction
In this paper, unless otherwise specified, $k$ always denotes an algebraically closed field with $\text{char}(k)=p>0$.
Let $C$ be a smooth projective curve over $k$. A classical question in theory of vector bundles over positive characteristic is whether every semistable vector bundle is strongly semistable or not. D. Gieseker [Gie73] constructed the first example of a semistable rank two vector bundle on a smooth projective curve over $k$ which is not strongly semistable.
In [LSZ13], Lan, Sheng and Zuo introduced the notion of strongly semistability of Higgs bundles, which generalized the notion of strongly semistability of vector bundle naturally, and proved that every rank two semistable Higgs bundle is strongly semistable ([LSZ13, Theorem 2.6]). Moreover, it was conjectured by Lan-Sheng-Zuo that every semistable Higgs bundle, therefore particularly every semistable bundle, is strongly semistable. The precise statement is given as follows:
Conjecture 1.1 ([LSZ13, Conjecture 1.5]).
Every semistable nilpotent [^1] The nilpotency condition is a necessary condition for a Higgs bundle to initiate a Higgs-de Rham flow. Higgs bundle of exponent $\leq p-1$ is strongly semistable.
In ([Li14, Theorem 1.1]), Li proved that every rank three semistable Higgs bundle is strongly semistable if $\text{char}(k)\geq 3$. A decisive answer to Conjecture 1.1 is the following result.
Theorem 1.1 ([LSZ17, Theorem A.1][Lan14, Theorem 5.5]).
Let $C$ be a smooth projective curve over $k$. Then every semistable Higgs bundle on $C$ with rank $\leq p$ is strongly semistable.
When the rank of a vector bundle over $k$ is greater than $p$, it is said to be big. It is not much known about the Conjecture 1.1 for semistable Higgs bundles of big rank, and the main results of this paper are summarized as follows:
Theorem 1.2.
Let $p$ be any prime number and $C$ be a smooth projective curve over $k$ with $\text{char}(k)=p$. The following two statements hold:
  1. (i)
    If $g(C)\leq 1$, then Conjecture 1.1 holds true;
  2. (ii)
    If $g(C)\geq 2$, then for any big rank $r$, there exists a rank $r$ semistable Higgs bundle over $C$ which are not strongly semistable. In particular, Conjecture 1.1 fails for some Higgs bundles of big rank.
Remark 1.1.
Note that for $C=\mathbb{P}^{1}$, Conjecture 1.1 holds trivially, as semistable Higgs bundles on $\mathbb{P}^{1}$ are direct sums of line bundles of the same degree, equipped with the zero Higgs field.
Organization of the paper
In the first part, we prove that the Lan-Sheng-Zuo conjecture holds on the elliptic curves (Proposition 3.2.1).
In the second part, firstly we show the existence of counter-examples of the Lan-Sheng-Zuo conjecture on certain varieties of general type with dimension $n\geq 2$ (Corollary 4.2.1), which is an immediate consequence of results in [Zho16]. By restriction theorem, this implies the existence of counter-examples for the Lan-Sheng-Zuo conjecture on generic curves in some linear system. After that, we provide an explicit construction of counter-examples of Lan-Sheng-Zuo conjecture for all possible cases on smooth projective curve (Proposition 4.2.1, Remark 4.2.3). In particular, an immediate consequence is that the assumption in Theorem 1.1 is sharp.
In the end of second part, we prove the existence of uniformizing flat bundle in the positive characteristic case (Lemma 4.3.1) and use it to provide a tensor product type of counter-examples of the Lan-Sheng-Zuo conjecture, which also provides an example to show the category of strongly semistable Higgs bundles is not a tensor category.
In the last part, we collect some partial results of the Lan-Sheng-Zuo conjecture on higher dimensional varieties.
Acknowledgement
The authors would like to express gratitude to Jianping Wang, who motivated the proof of Lemma 4.3.1, and his comments on higher dimensional case, which improve the Proposition 5.1. Also, the authors would like to thank Lingguang Li, for his careful reading of the manuscript and for providing invaluable comments that improved the paper.
The second named author is supported by the Chinese Academy of Sciences Project for Young Scientists in Basic Research (Grant No. YSBR-032).
2. Preliminaries
Let $X$ be an $n$-dimensional smooth projective variety over $k$.
2.1. Slope stability of vector bundle
Let $\omega$ be an ample divisor on $X$.
Definition 2.1.1.
Let $\mathcal{E}$ be a torsion-free bundle over $X$. The $\omega$-slope of $\mathcal{E}$ is defined as
$\mu_{\omega}(\mathcal{E})=\frac{c_{1}(\mathcal{E})\cdot\omega^{n-1}}{\operatorname{rk}\mathcal{E}}.$
Definition 2.1.2.
A torsion-free sheaf $\mathcal{E}$ on $X$ is called $\omega$-semistable (resp. stable), if $\mu_{\omega}(\mathcal{F})\leq\mu_{\omega}(\mathcal{E})$ (resp. $\mu_{\omega}(\mathcal{F})<\mu_{\omega}(\mathcal{E})$) for all subsheaves $\mathcal{F}$ of $\mathcal{E}$.
Remark 2.1.1.
In the curve case, it suffices to test all subbundles $\mathcal{F}$ of $\mathcal{E}$.
Theorem 2.1.1 ([HN74]).
Let $\mathcal{E}$ be a torsion-free sheaf on $X$. Then there exists a unique filtration $\Sigma_{\omega}$,
$0=\mathcal{E}_{0}\subsetneq\mathcal{E}_{1}\subsetneq\dots\subsetneq\mathcal{E}_{s}=\mathcal{E},$
which is called the Harder-Narasimhan filtration, such that
  1. $(1)$
    $\operatorname{Gr}_{i}(\Sigma_{\omega})=\mathcal{E}_{i}/\mathcal{E}_{i+1}$ is a torsion-free $\omega$-semistable sheaf for each $i$;
  2. $(2)$
    $\mu_{\omega}(\operatorname{Gr}_{i}(\Sigma_{\omega}))$ is a strictly decreasing function in $i$.
Sketch of existence.
Put $\mu_{\omega}^{\max}(\mathcal{E}):=\sup\{\mu_{\omega}(\mathcal{F})\mid 0\neq\mathcal{F}\subseteq\mathcal{E}\text{ a coherent subsheaf}\}$. Then we need to prove that
  1. $(1)$
    $\mu_{\omega}^{\max}(\mathcal{E})<\infty$;
  2. $(2)$
    There exists a saturated subsheaf $\mathcal{F}_{1}\subseteq\mathcal{E}$ with maximal slope.
After that, suppose both $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$ are coherent subsheaves of rank $r_{1}$ and $r_{2}$ with maximal slope. By the following exact sequence
$0\to\mathcal{F}_{1}\cap\mathcal{F}_{2}\to\mathcal{F}_{1}\oplus\mathcal{F}_{2}\to\mathcal{F}_{1}+\mathcal{F}_{2}\to 0,$
one has
$c_{1}(\mathcal{F}_{1}+\mathcal{F}_{2})$ $=c_{1}(\mathcal{F}_{1})+c_{1}(\mathcal{F}_{2})-c_{1}(\mathcal{F}_{1}\cap\mathcal{F}_{2})$
$\operatorname{rk}(\mathcal{F}_{1}+\mathcal{F}_{2})$ $=\operatorname{rk}(\mathcal{F}_{1})+\operatorname{rk}(\mathcal{F}_{2})-\operatorname{rk}(\mathcal{F}_{1}\cap\mathcal{F}_{2}).$
Then
$\operatorname{rk}(\mathcal{F}_{1}+\mathcal{F}_{2})\mu_{\omega}(\mathcal{F}_{1}+\mathcal{F}_{2})$ $=r_{1}\mu_{\omega}(\mathcal{F}_{1})+r_{2}\mu_{\omega}(\mathcal{F}_{2})-\operatorname{rk}(\mathcal{F}_{1}\cap\mathcal{F}_{2})\mu_{\omega}(\mathcal{F}_{1}\cap\mathcal{F}_{2})$
$\geq(r_{1}+r_{2})\mu_{\omega}^{\max}(\mathcal{E})-\operatorname{rk}(\mathcal{F}_{1}\cap\mathcal{F}_{2})\mu_{\omega}^{\max}(\mathcal{E})$
$=\operatorname{rk}(\mathcal{F}_{1}+\mathcal{F}_{2})\mu_{\omega}^{\max}(\mathcal{E}).$
This shows $\mathcal{F}_{1}+\mathcal{F}_{2}$ also has the maximal slope. By adding all these subsheaves together, this gives the maximal $\omega$-destabilizing subsheaf $\mathcal{E}_{1}$. We repeat above process to obtain the maximal $\omega$-destabilizing subsheaf of $\mathcal{E}/\mathcal{E}_{1}$, and consider its preimage to obtain $\mathcal{E}_{2}$, that is, $\mathcal{E}_{2}/\mathcal{E}_{1}$ is the maximal destabilizer of $\mathcal{E}/\mathcal{E}_{1})$. The slope function satisfies the decreasing condition $\mu_{\omega}(\mathcal{E}_{1})>\mu_{\omega}(\mathcal{E}_{2}/\mathcal{E}_{1})$, otherwise we would have $\mu_{\omega}(\mathcal{E}_{1})\leq\mu_{\omega}(\mathcal{E}_{2})$, a contradiction. ∎
Remark 2.1.2.
The maximal $\omega$-destabilizing subsheaf of $\mathcal{E}$ is characterized by the following properties:
  1. $(1)$
    $\mu_{\omega}(\mathcal{E}_{1})\geq\mu_{\omega}(\mathcal{F})$ for every coherent subsheaf $\mathcal{F}$ of $\mathcal{E}$;
  2. $(2)$
    If $\mu_{\omega}(\mathcal{E}_{1})=\mu_{\omega}(\mathcal{F})$ for $\mathcal{F}\subset\mathcal{E}$, then $\mathcal{F}\subset\mathcal{E}_{1}$.
Remark 2.1.3.
The $\omega$-semistable filtration of the dual sheaf $\mathcal{E}^{*}$ is essentially the same as that of $\mathcal{E}$, with each entry substituted by the duals of the quotient $\mathcal{E}/\mathcal{E}_{s-i}$.
Lemma 2.1.1 ([Miy87, Proposition 3.2]).
Let $f$ be a separable surjective $k$-morphism of a smooth complete curve $C^{\prime}$ onto $C$. Then a vector bundle $\mathcal{E}$ is semistable if and only if $f^{*}\mathcal{E}$ is semistable.
Definition 2.1.3.
Let $X$ be a smooth projective variety over $k$ with $\text{char}(k)=p>0$ and $F$ be the absolute Frobenius map on $X$. A vector bundle $\mathcal{E}$ on $X$ is called strongly semistable, if $(F^{*})^{n}\mathcal{E}$ is semistable for all $n\in\operatorname{\mathbb{Z}}_{\geq 0}$.
2.2. Higgs-de Rham flow
Let $X$ be a smooth projective variety over $k$ with $\text{char}(k)=p>0$.
Definition 2.2.1.
A Higgs bundle on $X$ is a pair $(\mathcal{E},\theta)$, where $\mathcal{E}$ is a vector bundle on $X$ and $\theta\colon\mathcal{E}\to\mathcal{E}\otimes\Omega_{X}^{1}$ is an $\mathcal{O}_{X}$-module morphism.
Definition 2.2.2.
A Higgs bundle $(\mathcal{E},\theta)$ is called $\omega$-semistable (resp. stable), if $\mu_{\omega}(\mathcal{F})\leq\mu_{\omega}(\mathcal{E})$ (resp. $\mu_{\omega}(\mathcal{F})<\mu_{\omega}(\mathcal{E})$) for all $\theta$-invariant subsheaves $\mathcal{F}$ of $\mathcal{E}$.
Definition 2.2.3.
A flat bundle over $X$ is a pair $(\mathcal{V},\nabla)$, where $\mathcal{V}$ is a vector bundle and $\nabla$ is an $k$-connection $\nabla\colon\mathcal{V}\to\mathcal{V}\otimes\Omega_{X}^{1}$ such that $\nabla^{2}=0$.
Definition 2.2.4.
A flat bundle $(\mathcal{V},\nabla)$ is called $\nabla$-semistable (resp. stable) with respect to $\omega$, if $\mu_{\omega}(\mathcal{F})\leq\mu_{\omega}(\mathcal{V})$ (resp. $\mu_{\omega}(\mathcal{F})<\mu_{\omega}(\mathcal{V})$) for all $\nabla$-invariant subsheaves $\mathcal{F}$ of $\mathcal{V}$.
Definition 2.2.5.
Let $(\mathcal{E},\theta)$ be a Higgs bundle on $X$.
  1. $(1)$
    We say that $(\mathcal{E},\theta)$ is nilpotent of exponent $<\ell$ if $(\mathcal{E},\theta)$ satisfies one of the following equivalent conditions:
    1. $(a)$
      There exists a filtration $0=\mathcal{F}^{m}\subset\mathcal{F}^{m-1}\subset\ldots\subset\mathcal{F}^{0}=(\mathcal{E},\theta)$ of length $m\leq\ell$ such that the associated graded Higgs bundle has zero Higgs field.
    2. $(b)$
      For any open subset $U\subset X$ and any collection $\left\{\partial_{x_{1}},\ldots,\partial_{x_{\ell}}\right\}$ of sections of $\mathcal{T}_{X}(U)$ we have $\theta\left(\partial_{x_{1}}\right)\ldots\theta\left(\partial_{x_{\ell}}\right)=0$.
  2. $(2)$
    We say that $(\mathcal{E},\theta)$ is nilpotent of exponent $=\ell$ if it is nilpotent of exponent $<\ell+1$ but not nilpotent of exponent $<\ell$ (for $\ell=0$ we require simply that $(\mathcal{E},\theta)$ is nilpotent of exponent $<1$).
Definition 2.2.6.
Let $(\mathcal{V},\nabla)$ be a flat bundle on $X$.
  1. $(1)$
    We say that the $p$-curvature $\psi$ of $(\mathcal{V},\nabla)$ is nilpotent of exponent $<\ell$ if $(\mathcal{V},\nabla)$ satisfies one of the following equivalent conditions:
    1. $(a)$
      There exists a filtration $0=\mathcal{F}^{m}\subset\mathcal{F}^{m-1}\subset\ldots\subset\mathcal{F}^{0}=(\mathcal{V},\nabla)$ of length $m\leq\ell$ such that the associated graded flat bundle has $p$-curvature zero.
    2. $(b)$
      For any open subset $U\subset X$ and any collection $\left\{\partial_{x_{1}},\ldots,\partial_{x_{\ell}}\right\}$ of sections of $\mathcal{T}_{X}(U)$ we have $\psi\left(\partial_{x_{1}}\right)\ldots\psi\left(\partial_{x_{\ell}}\right)=0$.
  2. $(2)$
    We say that the $p$-curvature of $(\mathcal{V},\nabla)$ is nilpotent of exponent $=\ell$ if it is nilpotent of exponent $<\ell+1$ but not nilpotent of exponent $<\ell$ (for $\ell=0$ we require simply that the $p$-curvature is nilpotent of exponent $<1$).
Notation 2.2.1.
The set consisting of Higgs bundles which are nilpotent of exponent $\leq p-1$ is denoted by $\operatorname{HIG}_{p-1}(X)$, and the set consisting of flat bundles which are nilpotent of exponent $\leq p-1$ is denoted by $\operatorname{MIC}_{p-1}(X)$.
A. Ogus and V. Vologodsky established the non-abelian Hodge theory in positive characteristic case, and a simple version can be stated as follows:
Theorem 2.2.1 ([OV07, Theorem 2.8]).
If $X$ is $W_{2}(k)$-liftable, then there is an equivalence of categories between $\operatorname{HIG}_{p-1}(X)$ and $\operatorname{MIC}_{p-1}(X)$, where the correspondence is given by the inverse Cartier transform.
Remark 2.2.1.
The inverse Cartier of a vector bundle $\mathcal{E}$ is the Frobenius pullback $F^{*}\mathcal{E}$. In the non-trivial Higgs field case, the inverse Cartier transform can be explained as an exponential twisting of Frobenius map by Higgs field in [LSZ15].
Definition 2.2.7.
Let $(\mathcal{E},\theta)$ be a Higgs bundle on $X$. A Higgs-de Rham flow with initial term $(\mathcal{E},\theta)$ is an infinite sequence
[…]
in which $C^{-1}$ is the inverse Cartier transform with respect to $W_{2}(k)$-lifting $X\hookrightarrow\widetilde{X}$ and $\mathcal{N}_{i}^{\bullet}$ is a Griffiths transverse filtration of $(\mathcal{V}_{i},\nabla_{i})$ such that $(\mathcal{E}_{i+1},\theta_{i+1})$ is the associated graded Higgs bundle for each $i$.
Definition 2.2.8.
A Higgs-de Rham flow with initial term $(\mathcal{E},\theta)$ is called $f$-periodic, if there exists a natural number $f>0$ such that there is an isomorphism of Higgs bundles between $(\mathcal{E},\theta)$ and the $f$-th term $(\mathcal{E}_{f},\theta_{f})$.
Definition 2.2.9.
A Higgs-de Rham flow is called a semistable Higgs-de Rham flow, if each $(\mathcal{E}_{i},\theta_{i})$ is a semistable Higgs bundle.
Definition 2.2.10.
A semistable Higgs bundle is called strongly semistable, if it initiates a semistable Higgs-de Rham flow.
Example 2.2.1.
If $\mathcal{E}$ is a strongly semistable vector bundle, then $(\mathcal{E},0)$ is a strongly semistable Higgs bundle since it initiates the following semistable Higgs-de Rham flow
[…]
where $\mathcal{N}_{i}^{\bullet}$ is the trivial filtration for each $i$. This shows the notion of strongly semistability for a Higgs bundle generalizes naturally the notion of strongly semistability for a vector bundle.
3. Lan-Sheng-Zuo conjecture on elliptic curves
3.1. Gr-polystable filtration
3.1.1. Motivation and first examples
In [Lan14], A. Langer claimed that every Simpson filtration of a flat bundle $(\mathcal{V},\nabla)$ can be refined such that the associated graded Higgs bundle is polystable. However, we point out that not every Simpson filtration can be refined to be a gr-polystable filtration.
Example 3.1.1.
Let $E$ be an ordinary elliptic curve over $k$. Consider the trivial extension
$0\to\mathcal{O}_{E}\to\mathcal{E}\to\mathcal{O}_{E}\to 0$
and Higgs field $\theta$ on $\mathcal{E}$ is given by the isomorphism $\mathcal{O}_{E}\to\mathcal{O}_{E}$. In this case, if we choose the canonical $W_{2}(k)$-lifting of $E$, then the flat bundle $(\mathcal{V},\nabla)=C^{-1}(\mathcal{E},\theta)$ is a non-trivial extension of flat bundles as follows:
$0\to(\mathcal{L}_{1},\nabla_{\operatorname{can}})\to(\mathcal{V},\nabla)\to(\mathcal{L}_{2},\nabla_{\operatorname{can}})\to 0,$
where $\mathcal{L}_{1}\cong\mathcal{L}_{2}\cong\mathcal{O}_{E}$. By [KS24, Lemma 2.2] we have that $\mathcal{V}\cong\mathcal{L}_{1}\oplus\mathcal{L}_{2}$ since by the choice of lifting the obstruction class for Frobenius lifting is zero. Then $0\subseteq\mathcal{L}_{2}\subseteq\mathcal{V}$ is a Simpson filtration with $\nabla(\mathcal{L}_{2})\subsetneq\mathcal{L}_{2}$, and thus the associated graded Higgs bundle is $(\mathcal{O}_{E}\oplus\mathcal{O}_{E},\theta)$, where Higgs field is given by the isomorphism $\theta\colon\mathcal{O}_{E}\to\mathcal{O}_{E}$. In other words, $(\mathcal{O}_{E}\oplus\mathcal{O}_{E},\theta)$ is one periodic.
Since $(\mathcal{O}_{E}\oplus\mathcal{O}_{E},\theta)$ is not polystable and the Simpson filtration $0\subseteq\mathcal{L}\subseteq\mathcal{V}$ cannot be refined anymore, this provides a counter-example of Langer’s claim. However, if we choose the trivial filtration, which is also a Simpson filtration, then the associated graded Higgs bundle $(\mathcal{O}_{E}\oplus\mathcal{O}_{E},0)$ is polystable.
Instead of considering refinement for an arbitrary Simpson filtration, we should prove that there always exists a Simpson filtration such that the associated graded Higgs bundle is polystable, which is called a gr-polystable filtration for convenience. For smooth projective curve of genus zero and one, it is easy to figure out which one is gr-polystable filtration.
Example 3.1.2.
The existence of gr-polystable filtration is trivial on $\mathbb{P}^{1}$ since every semistable Higgs bundle on $\mathbb{P}^{1}$ is polystable.
Example 3.1.3.
Let $E$ be an elliptic curve and $(\mathcal{V},\nabla)$ be $\nabla$-semistable flat bundle on $E$. Since every $\nabla$-semistable flat bundle on an elliptic curve is semistable as a vector bundle, so $\mathcal{V}$ is a semistable vector bundle. Thus an obvious Simpson filtration of $(\mathcal{V},\nabla)$ is the trivial filtration, and the gr-polystable filtration is given by the Jordan-Hölder filtration of $\mathcal{V}$.
3.2. Lan-Sheng-Zuo conjecture on elliptic curves
In this section, we provide two different proofs of Lan-Sheng-Zuo conjecture on an elliptic curve $E$. One of the proofs is based on the existence of gr-polystable filtration and Lemma 3.2.1, and another proof is based on the fact that every semistable bundle on an elliptic curve is strongly semistable.
Lemma 3.2.1 ([FGPN13, Proposition 4.3]).
Let $(\mathcal{E},\theta)$ be a stable Higgs bundle on $E$. Then $\mathcal{E}$ is stable as a vector bundle.
Proposition 3.2.1.
Let $E$ be an elliptic curve over $k$. Then every semistable Higgs bundle $(\mathcal{E},\theta)$ which is nilpotent of exponent $\leq p-1$ is strongly semistable.
First proof.
Suppose $(\mathcal{V},\nabla)=C^{-1}(\mathcal{E},\theta)$ is the flat bundle obtained by inverse Cartier transform. Choose the gr-polystable filtration on $(\mathcal{V},\nabla)$ and $(\mathcal{E}^{\prime},\theta^{\prime})$ denotes the associated graded Higgs bundle. Now we claim that $\theta^{\prime}=0$. Indeed, since $(\mathcal{E}^{\prime},\theta^{\prime})$ is polystable, we write it as
$(\mathcal{E}^{\prime},\theta^{\prime})=\bigoplus_{i=1}^{n}(\mathcal{E}_{i},\theta_{i}),$
where $(\mathcal{E}_{i},\theta_{i})$ are stable Higgs bundles. By Lemma 3.2.1 we have $\mathcal{E}_{i}$ is stable as vector bundle for all $i$, and thus $\theta_{i}$ is a scalar endomorphism. Since $\theta_{i}$ is nilpotent, we must have $\theta_{i}=0$ for all $i$. ∎
Second proof.
Let $(\mathcal{V},\nabla)=C^{-1}(\mathcal{E},\theta)$. Since $\mathcal{V}$ is semistable as a vector bundle, we may choose trivial filtration on $(\mathcal{V},\nabla)$ as its Simpson filtration, then the associated graded Higgs bundle is $(\mathcal{V},0)$. Since every semistable bundle on an elliptic curve is strongly semistable, $(\mathcal{E},\theta)$ initiates a semistable Higgs de Rham flow as follows
[…]
where $\mathcal{N}_{i}^{\bullet}$ is the trivial filtration for each $i$. ∎
4. Counter-examples of Lan-Sheng-Zuo conjecture on higher genus curves
In this section, we will provide the construction of counter-example of Lan-Sheng-Zuo conjecture, and the construction is mainly based on the Frobenius destabilized bundle constructed by the direct images of Frobenius map.
4.1. Stability of direct images of absolute Frobenius map
Let $f\colon X\to Y$ be a finite map between smooth projective curves of genus $\geq 2$. In an unpublished note [Bea00], Beauville conjectured that $f_{*}\mathcal{E}$ is stable for a general vector bundle $\mathcal{E}$ and he proved several cases. A generic version of this conjecture is proved in [CLV24] recently.
4.1.1. Frobenius destabilized bundle
In this section, we introduce several results about Beauville’s conjecture when $f$ is the absolute Frobenius map, and this provides a lot of semistable bundle which is not strongly semistable.
However, the following result shows that the Frobenius destabilized bundles only exist on general type varieties.
Theorem 4.1.1 ([MR83, Theorem 2.1], [Sun10, Corollary 3.10]).
Let $X$ be a smooth projective variety of dimension $n$ over $k$.
  1. $(1)$
    If $\mu_{\max}(\Omega_{X}^{1})\leq 0$, then all semistable sheaves on $X$ are strongly semistable.
  2. $(2)$
    If $\mu_{\max}(\Omega_{X}^{1})<0$, then all stable sheaves on $X$ are strongly stable.
Lemma 4.1.1.
Let $X$ be a smooth projective variety of dimension $n$ over $k$ and $F$ be the absolute Frobenius map. Let $\mathcal{E}$ be a vector bundle of rank $r$. Then
  1. $(1)$
    $\operatorname{rk}(F_{*}\mathcal{E})=p^{n}r$ and
    $c_{1}(F_{*}\mathcal{E})=\frac{r(p^{n}-p^{n-1})}{2}K_{X}+p^{n-1}c_{1}(\mathcal{E}).$
  2. $(2)$
    $\operatorname{rk}(F^{*}\mathcal{E})=r$ and $c_{1}(F^{*}\mathcal{E})=pc_{1}(\mathcal{E})$.
Proof.
It follows from the Grothendieck-Riemann-Roch theorem. ∎
Corollary 4.1.1.
Let $C$ be a smooth projective curve of genus $g$ over $k$ and $F$ be the absolute Frobenius map. Let $\mathcal{E}$ be a vector bundle of degree $d$ and rank $r$. Then
  1. $(1)$
    $\operatorname{rk}(F_{*}\mathcal{E})=pr$ and $\deg(F_{*}\mathcal{E})=r(p-1)(g-1)+d$.
  2. $(2)$
    $\operatorname{rk}(F^{*}\mathcal{E})=r$ and $\deg(F^{*}\mathcal{E})=pd$.
Theorem 4.1.2 ([Sun08, Theorem 2.2]).
Let $X$ be a smooth projective curve of genus $g\geq 1$. Then $F_{*}\mathcal{E}$ is semistable whenever $\mathcal{E}$ is semistable. If $g\geq 2$, then $F_{*}\mathcal{E}$ is stable whenever $\mathcal{E}$ is stable.
Proposition 4.1.1.
Let $C$ be a smooth projective curve of genus $g\geq 2$ over $k$ and $\mathcal{E}$ be a vector bundle. Then the adjunction map $F^{*}F_{*}\mathcal{E}\to\mathcal{E}$ is surjective and
$\mu(F^{*}F_{*}\mathcal{E})=\mu(\mathcal{E})+(p-1)(g-1)>\mu(\mathcal{E}).$
In particular, $F^{*}F_{*}\mathcal{E}$ is unstable.
Proof.
The surjectivity of the adjunction map can be checked by a local calculation and the formula for the slope follows from Corollary 4.1.1. ∎
This shows one can obtain lots of examples of Frobenius destabilized bundle by considering the direct image of Frobenius map, and the Harder-Narasimhan filtration of this kind of examples is given by its canonical filtration as introduced in the following section.
4.1.2. Canonical filtration
Let $X$ be a smooth projective variety of dimension $n$ over $k$ and $F$ be the absolute Frobenius map.
Definition 4.1.1.
Let $\mathcal{E}$ be a vector bundle on $X$. Then the canonical filtration on $\mathcal{V}=F^{*}F_{*}\mathcal{E}$ is a decreasing filtration
$0=\mathcal{V}_{n(p-1)+1}\subset\mathcal{V}_{n(p-1)}\subset\dots\subset\mathcal{V}_{0}=\mathcal{V},$
where $\mathcal{V}_{1}=\ker\{\mathcal{V}=F^{*}F_{*}\mathcal{E}\twoheadrightarrow\mathcal{E}\}$ and
$\mathcal{V}_{\ell+1}=\ker\{\mathcal{V}_{\ell}\stackrel{{\scriptstyle\nabla}}{{\longrightarrow}}\mathcal{V}\otimes\Omega_{X}^{1}\to\mathcal{V}/\mathcal{V}_{\ell}\otimes\Omega_{X}^{1}\}.$
Lemma 4.1.2 ([JRXY06, Theorem 5.3]).
Let $C$ be a smooth projective curve of genus $g$ over $k$ and $\mathcal{E}$ be a vector bundle on $C$. Then the following properties hold for canonical filtration of $F^{*}F_{*}\mathcal{E}$:
  1. $(1)$
    $\mathcal{V}_{0}/\mathcal{V}_{1}\cong\mathcal{E},\nabla(\mathcal{V}_{\ell+1})\subseteq\mathcal{V}_{\ell}\otimes\Omega_{C}^{1}$ for $\ell\geq 1$.
  2. $(2)$
    $\mathcal{V}_{\ell}/\mathcal{V}_{\ell+1}\stackrel{{\scriptstyle\nabla}}{{\longrightarrow}}(\mathcal{V}_{\ell-1}/\mathcal{V}_{\ell})\otimes\Omega_{C}^{1}$ is an isomorphism for $11\leq\ell\leq p-$ 1.
  3. $(3)$
    If $g\geq 2$ and $\mathcal{E}$ is semistable, then the canonical filtration is nothing but the Harder-Narasimhan filtration of $F^{*}F_{*}\mathcal{E}$.
Theorem 4.1.3 ([Sun08, Theorem 3.7]).
Let $X$ be a smooth projective variety of dimension $n$ over $k$ and $\mathcal{E}$ be a vector bundle on $X$. Then the following properties hold for canonical filtration of $F^{*}F_{*}\mathcal{E}$:
  1. $(1)$
    $\mathcal{V}_{0}/\mathcal{V}_{1}\cong\mathcal{E},\nabla(\mathcal{V}_{\ell+1})\subseteq\mathcal{V}_{\ell}\otimes\Omega_{X}^{1}$ for $\ell\geq 1$.
  2. $(2)$
    $\mathcal{V}_{\ell}/\mathcal{V}_{\ell+1}\xrightarrow{\nabla}(\mathcal{V}_{\ell-1}/\mathcal{V}_{\ell})\otimes\Omega_{X}^{1}$ are injective morphisms of vector bundles for $1\leq\ell\leq n(p-1)$, which induced isomorphisms
    $\nabla^{\ell}\colon\mathcal{V}_{\ell}/\mathcal{V}_{\ell+1}\cong\mathcal{E}\otimes_{\mathcal{O}_{X}}\mathrm{T}^{\ell}(\Omega_{X}^{1}),\quad 0\leq\ell\leq n(p-1).$
    The vector bundle $\mathrm{T}^{\ell}(\Omega_{X}^{1})$ is suited in the exact sequence
    00 0 $\rightarrow\operatorname{Sym}^{\ell-\ell(p)\cdot p}(\Omega_{X}^{1})\otimes F^{*}\Omega_{X}^{\ell(p)}\xrightarrow{\phi}\operatorname{Sym}^{\ell-(\ell(p)-1)\cdot p}(\Omega_{X}^{1})\otimes F^{*}\Omega_{X}^{\ell(p)-1}$
    $\rightarrow\cdots\rightarrow\operatorname{Sym}^{\ell-q\cdot p}(\Omega_{X}^{1})\otimes F^{*}\Omega_{X}^{q}\xrightarrow{\phi}\operatorname{Sym}^{\ell-(q-1)\cdot p}(\Omega_{X}^{1})\otimes F^{*}\Omega_{X}^{q-1}$
    $\rightarrow\cdots\rightarrow\operatorname{Sym}^{\ell-p}(\Omega_{X}^{1})\otimes F^{*}\Omega_{X}^{1}\xrightarrow{\phi}\operatorname{Sym}^{\ell}(\Omega_{X}^{1})\rightarrow\mathrm{T}^{\ell}(\Omega_{X}^{1})\rightarrow 0$
    where $\ell(p)\geq 0$ is the integer such that $\ell-\ell(p)\cdot p<p$.
Lemma 4.1.3 ([Zho16, Lemma 2.4]).
Let $(X,\omega)$ be a polarized smooth projective variety over $k$ such that $\Omega_{X}^{1}$ is strongly $\omega$-semistable and $\mu_{\omega}(\Omega_{X}^{1})>0$. Let $\mathcal{E}$ be a strongly $\omega$-semistable bundle. Then $F_{*}\mathcal{E}$ is $\omega$-semistable and the canonical filtration of $F^{*}F_{*}\mathcal{E}$ coincides with the Harder-Narasimhan filtration.
4.2. Counter-example of Lan-Sheng-Zuo conjecture on higher genus curve
This section is divided into three parts, firstly we point out there always exists counter-example of Lan-Sheng-Zuo conjecture on some higher dimensional varieties of general type, which is an immediate consequence of Lemma 4.1.3. As a consequence, this proves that the generic existence of counter-examples of Lan-Sheng-Zuo conjecture.
In the second part, we give a more explicit way to construct counter-examples of all possible curve case, that is, given any $g\in\operatorname{\mathbb{Z}}_{\geq 2}$, prime $p>0$ and $r\in\operatorname{\mathbb{Z}}_{>p}$, for any smooth projective curve of genus $g$ over $k$ with $\text{char}(k)=p$, there exists a counter-example of rank $r$ for Lan-Sheng-Zuo conjecture. As a consequence, this implies that the result in Theorem 1.1 is sharp.
In the third part, we prove the existence of uniformizing flat bundle in the positive characteristic case and use it to provide a counter-example which is constructed by the tensor product of two strongly semistable bundles. This also provides an explicit counter-example to show the category of strongly semistable Higgs bundles is not a tensor category.
4.2.1. Existence of counter-example of Lan-Sheng-Zuo conjecture on some general type varieties of dimension $n\geq 2$
Let $(X,\omega)$ be a polarized smooth projective variety of dimension $n\geq 2$ such that $\Omega_{X}^{1}$ is strongly $\omega$-semistable and $\mu_{\omega}(\Omega_{X}^{1})>0$ and $F$ be the absolute Frobenius map on $X$. For example, we can choose $X=C\times\dots\times C$, where $C$ is a smooth projective curve of general type and $\omega=\mathcal{O}_{C}(1)\boxtimes\dots\boxtimes\mathcal{O}_{C}(1)$.
Let $\mathcal{L}$ be a line bundle on $X$ and $\mathcal{E}$ be the direct image of $\mathcal{L}$ under $F$. Then by Lemma 4.1.3 we have $\mathcal{E}$ is $\omega$-semistable, and the Harder-Narasimhan filtration of $F^{*}\mathcal{E}$ is given by its canonical filtration
$0=\mathcal{V}_{n(p-1)+1}\subset\mathcal{V}_{n(p-1)}\subset\dots\subset\mathcal{V}_{0}=F^{*}\mathcal{E}.$
The canonical filtration satisfies the Griffiths transversality and the associated graded Higgs bundle $(\mathcal{E},\theta)=(\bigoplus_{i=1}^{n(p-1)+1}\mathcal{E}_{i},\bigoplus_{i=1}^{n(p-1)}\theta_{i})$ such that $\theta_{i}$ is injective for all $i=1,\dots,n(p-1)$.
By construction, we have
  1. (1)
    $(\mathcal{E},\theta)$ is a semistable Higgs bundle since the slopes of graded pieces in Harder-Narasimhan filtration are strictly decreasing;
  2. (2)
    $(\mathcal{E},\theta)$ is nilpotent of exponent $=n(p-1)$, which is larger than $p$ when $n\geq 2$.
This provides lots of counter-examples of Lan-Sheng-Zuo conjecture on higher dimensional smooth projective varieties of general type.
Corollary 4.2.1.
Let $(X,\omega)$ be a polarized smooth projective variety of dimension $n\geq 2$ with $\Omega_{X}^{1}$ strongly $\omega$-semistable and $\mu_{\omega}(\Omega_{X}^{1})>0$. Then Lan-Sheng-Zuo conjecture fails on $X$.
Remark 4.2.1.
The above argument cannot be generalized to arbitrary general type varieties, and thus it’s still not clear the existence of counter-examples of Lan-Sheng-Zuo conjecture on any general type varieties of dimension $n\geq 2$.
Remark 4.2.2.
According to the restriction theorem of Higgs bundles, the above example provides a curve case counter-example of Lan-Sheng-Zuo conjecture by restricting on generic curves, as the stability and exponent of nilpotency do not change when restricting to the generic open subset.
4.2.2. Counter-example of Lan-Sheng-Zuo conjecture on curves
Let $C$ be a smooth projective curve of genus $g\geq 2$ over $k$ with $\text{char}(k)=p>0$ and $F$ be the absolute Frobenius map on $C$.
Let $\mathcal{L}$ be a line bundle on $C$ and $\mathcal{E}_{1}=F_{*}\mathcal{L}$ be the Frobenius pushforward of $\mathcal{L}$. By Theorem 4.1.2 and Corollary 4.1.1 we have $\mathcal{E}_{1}$ is a stable rank $p$ bundle of degree $\deg\mathcal{L}+(p-1)(g-1)$. Moreover, by Lemma 4.1.2 the unique Simpson filtration on $(F^{*}\mathcal{E}_{1},\nabla_{\operatorname{can}})$ is given by the canonical filtration of $F^{*}\mathcal{E}_{1}$, and the associated graded Higgs bundle with respect to the canonical filtration is
$(\mathcal{O}_{C}\oplus K_{C}\oplus K_{C}^{\otimes 2}\oplus\dots\oplus K_{C}^{\otimes(p-1)},\bigoplus_{i=1}^{p-1}\theta_{i})\otimes(\mathcal{L},0),$
where $\theta_{i}\colon K_{C}^{\otimes i-1}\to K_{C}^{\otimes i}$ is an isomorphism for each $i=1,\dots,p-1$.
If we choose $\mathcal{L}$ to be the Frobenius pullback of a line bundle $\mathcal{L}_{2}$, then $\mathcal{L}\otimes K_{C}^{\otimes p}=F^{*}(\mathcal{L}_{2}\otimes K_{C})$ can also be equipped with a canonical connection $\nabla_{\operatorname{can}}$. Thus we define $\mathcal{V}=F^{*}\mathcal{E}_{1}\oplus\left(\mathcal{L}\otimes K_{C}^{\otimes p}\right)$ together with a flat connection $\nabla$ given by
$\begin{pmatrix}\nabla_{\operatorname{can}}&\theta\\ 0&\nabla_{\operatorname{can}}\end{pmatrix},$
where $\theta\colon\mathcal{L}\otimes K_{C}^{\otimes p}\to\mathcal{L}\otimes K_{C}^{\otimes p}\subseteq F^{*}\mathcal{E}_{1}\otimes K_{C}$ is an isomorphism.
It’s clear that the nilpotent of exponent of $p$-curvature of $(\mathcal{V},\nabla)$ is $\leq 1$, since $0\subset F^{*}\mathcal{E}_{1}\subset\mathcal{V}$ is a filtration of $\mathcal{V}$ such that the $p$-curvature of each graded piece is zero. Thus $(\mathcal{E},\theta)=C(\mathcal{V},\nabla)$ is well-defined. Moreover, $(\mathcal{E},\theta)$ is a stable Higgs bundle, since $(\mathcal{V},\nabla)$ is $\nabla$-stable.
Proposition 4.2.1.
The stable Higgs bundle $(\mathcal{E},\theta)$ provides a counter-example of rank $p+1$ of Lan-Sheng-Zuo conjecture on smooth projective curves of $g\geq 2$.
Proof.
It suffices to note that the Simpson filtration on $(\mathcal{V},\nabla)=C^{-1}(\mathcal{E},\theta)$ is given by
$0\subset\mathcal{L}\otimes K_{C}^{\otimes p}\subset\mathcal{L}\otimes K_{C}^{\otimes p}\oplus\mathcal{N}^{\bullet}\subset\mathcal{V},$
where $\mathcal{N}$ is the canonical filtration of $F^{*}\mathcal{E}_{1}$. Thus the Simpson filtration of $(\mathcal{V},\nabla)$ is unique and $(\mathcal{E},\theta)$ provides a counter-example of Lan-Sheng-Zuo conjecture since the associated graded Higgs bundle is
$(\mathcal{O}_{C}\oplus K_{C}\oplus K_{C}^{\otimes 2}\oplus\dots\oplus K_{C}^{\otimes p},\bigoplus_{i=1}^{p}\theta_{i})\otimes(\mathcal{L},0),$
which is a stable Higgs bundle with nilpotent of exponent $=p$. ∎
Remark 4.2.3.
The existence of a counter-example with rank $p+1$ can provide counter-examples with rank $r$ for any $r\in\operatorname{\mathbb{Z}}_{\geq p+1}$. Indeed, by considering the direct sum of a counter-example of rank $k$ with a line bundle with appropriate degree, it provides a counter-example of rank $k+1$.
4.3. Tensor product type counter-example of Lan-Sheng-Zuo conjecture on curves
In this section, we construct an example to show that the tensor product of two strongly semistable Higgs bundle may not be strongly semistable. This provides a tensor product type counter-example of rank $2p$ for Lan-Sheng-Zuo conjecture on smooth projective curve $C$ of genus $g\geq 2$, also shows the category consisting of strongly semistable Higgs bundles is not a tensor category.
4.3.1. The existence of uniformizing flat bundle and oper
The construction of the tensor product type counter-example is based on the so-called uniformizing flat bundle (Definition 4.3.1). In the complex number case, N. Hitchin introduced the uniformizing Higgs bundle $(\mathcal{E}_{\operatorname{unif}},\theta_{\operatorname{unif}}):=(K_{C}^{-\frac{1}{2}}\oplus K_{C}^{\frac{1}{2}},\theta)$, where $\theta\colon K_{C}^{\frac{1}{2}}\to K_{C}^{-\frac{1}{2}}\otimes K_{C}$ is the isomorphism, to recover the uniformization theorem of a compact Riemann surface by solving the Yang-Mills-Higgs equation on the uniformizing Higgs bundle in [Hit87]. An important observation is that the unique solution to the Yang-Mills-Higgs equation associates to the base curve a polarized $\operatorname{\mathbb{C}}$-variation of Hodge structure.
Definition 4.3.1.
A flat bundle $(\mathcal{V}_{\operatorname{unif}},\nabla_{\operatorname{unif}})$ together with a filtration is called a uniformizing flat bundle, if the associated graded Higgs bundle is the uniformizing Higgs bundle $(\mathcal{E}_{\operatorname{unif}},\theta_{\operatorname{unif}})$.
Let’s recall the following basic facts about the existence of flat connection which we will use in the proof of the existence of uniformizing flat bundle. In the complex number case, M. Atiyah proved that a holomorphic vector bundle $\mathcal{E}$ over curve admitting a flat connection if and only if the degree of each indecomposable component of $\mathcal{E}$ is zero in [Ati57]. The similar result holds in positive characteristic case.
Proposition 4.3.1 ([BS06, Proposition 3.1]).
Let $C$ be a smooth projective curve defined over $k$, where $k$ is algebraically closed with characteristic $p>0$. Let $\mathcal{E}$ be a vector bundle on $C$. Then it admits a connection if and only if the degree of each indecomposable component of $\mathcal{E}$ is a multiple of $p$.
Lemma 4.3.1.
Let $C$ be a smooth projective curve of genus $g\geq 2$ over $k$. Then uniformizing flat bundle $(\mathcal{V}_{\operatorname{unif}},\nabla_{\operatorname{unif}})$ always exists.
Proof.
If $p\mid g-1$, then by Proposition 4.3.1 there exists flat connections $\nabla_{+}$ and $\nabla_{-}$ on $K_{C}^{\frac{1}{2}}$ and $K_{C}^{-\frac{1}{2}}$ respectively. Then we define $\mathcal{V}_{\operatorname{unif}}=K_{C}^{\frac{1}{2}}\oplus K_{C}^{-\frac{1}{2}}$ and flat connection $\nabla_{\operatorname{unif}}$ is defined by
$\begin{pmatrix}\nabla_{+}&0\\ \theta&\nabla_{-}\end{pmatrix},$
where $\theta\colon K_{C}^{\frac{1}{2}}\to K_{C}^{-\frac{1}{2}}\otimes K_{C}$ is the isomorphism.
If $p\nmid g-1$, we choose $\mathcal{V}_{\operatorname{unif}}$ to be the unique non-trivial extension
$0\to K_{C}^{\frac{1}{2}}\to\mathcal{V}\to K_{C}^{-\frac{1}{2}}\to 0,$
which is given by a non-trivial element in $\operatorname{Ext}^{1}(K_{C}^{-\frac{1}{2}},K_{C}^{\frac{1}{2}})\cong H^{1}(C,K_{C})\cong k$. Then $\mathcal{V}$ is indecomposable of degree zero, and thus again by Proposition 4.3.1 there exists a flat connection $\nabla_{\operatorname{unif}}$ on $\mathcal{V}_{\operatorname{unif}}$. Moreover, the subbundle $K_{C}^{\frac{1}{2}}$ is not $\nabla$-invariant, otherwise $K_{C}^{\frac{1}{2}}$ admits a flat connection, which contradicts to $p\nmid g-1$ by Proposition 4.3.1. ∎
Remark 4.3.1.
The existence of uniformizing flat bundle over a generic smooth projective curve is a direct consequence of the periodicity of the uniformizing Higgs bundle, which is stated as follows:
Theorem 4.3.1 ([LSYZ19, Corollary 2.2]).
Let $C$ be a generic smooth projective curve of genus $g\geq 2$. Then there exists a $W_{2}(k)$-lifting $C\hookrightarrow\widetilde{C}$ such that the uniformizing Higgs bundle with respect to the $C\hookrightarrow\widetilde{C}$ is two-periodic. More precisely, there exists the following Higgs-de Rham flow
[…]
where $\mathcal{L}$ is a certain two-torsion line bundle.
This also shows uniformizing flat bundle is not unique, since the vector bundle $\mathcal{V}^{\prime}$ is a non-trivial extension of $K_{C}^{-\frac{p}{2}}$ by $K_{C}^{\frac{p}{2}}$, which is not isomorphic to uniformizing flat bundles constructed in Lemma 4.3.1.
The proof of the existence of uniformizing flat bundle can be generalized to higher ranks, which gives a proof of the existence of opers of degree zero in any rank.
Definition 4.3.2.
Let $X$ be a smooth projective variety. An oper on $X$ is a triple $(\mathcal{V},\nabla,\mathcal{F}^{\bullet})$ consists of a vector bundle $\mathcal{V}$ on $X$ together with a flat connection $\nabla$ and a Griffiths transversal filtration $\mathcal{F}^{\bullet}$ such that the quotients $\mathcal{F}^{i}/\mathcal{F}^{i+1}$ are line bundles on $X$ and the $\mathcal{O}_{X}$-linear homomorphisms $\theta^{i}\colon\mathcal{F}^{i}\to\mathcal{F}^{i+1}$ induced by the flat connection are isomorphisms.
Proposition 4.3.2.
Let $C$ be a smooth projective curve over $k$ with $\text{char}(k)=p>0$. Then for any $r\in\operatorname{\mathbb{Z}}_{>0}$, there exists an oper with rank $r$ and degree zero on $C$.
Proof.
The $r=1$ case is trivial and $r=2$ case is exactly the existence of uniformizing flat bundle. If $p\mid g-1$, then the construction is the same as the rank two case, so without lose of generality we may assume $p\nmid g-1$.
Under this assumption, we will prove there exists an oper $(\mathcal{V},\nabla,\mathcal{F}^{\bullet})$ of rank $r$ and degree zero such that $\mathcal{V}$ is an indecomposable vector bundle by induction on rank $r$.
Suppose the statement holds for the cases rank $\leq r-1$.
  1. $(1)$
    If $r$ is odd, then by induction hypothesis there exists an oper $(\mathcal{V}_{r-1},\nabla_{r-1},\mathcal{F}_{r-1}^{\bullet})$ such that $\mathcal{V}_{r-1}$ is indecomposable and the underlying vector bundle of the associated Higgs bundle is
    $K_{C}^{-\frac{(r-2)}{2}}\oplus\dots\oplus K_{C}^{-\frac{1}{2}}\oplus K_{C}^{\frac{1}{2}}\oplus\dots\oplus K_{C}^{\frac{r-2}{2}}.$
    If $H^{1}(C,K_{C}^{\frac{r-1}{2}}\otimes\mathcal{V}_{r-1}\otimes K_{C}^{\frac{1}{2}})\neq 0$, then there exists a non-trivial extension
    $0\to\mathcal{V}_{r-1}\otimes K_{C}^{\frac{1}{2}}\to\mathcal{V}_{r}\to K_{C}^{-\frac{(r-1)}{2}}\to 0$
    such that $\mathcal{V}_{r}$ is of rank $r$ and degree zero. Furthermore, the vector bundle $\mathcal{V}_{r}$ is indecomposable together with a filtration $\mathcal{F}_{r}^{\bullet}$ given by
    $0\subset\mathcal{F}_{r-1}^{\bullet}\otimes K_{C}^{\frac{1}{2}}\subset\mathcal{V}_{r}.$
    By Proposition 4.3.1, there exists a flat connection $\nabla_{r}$ on $\mathcal{V}_{r}$, and $\mathcal{F}_{r}^{\bullet}$ is a Griffiths transverse filtration with respect to $\nabla_{r}$, since $p\nmid g-1$. This shows $(\mathcal{V}_{r},\nabla_{r},\mathcal{F}^{\bullet}_{r})$ is a desired oper. Now it suffices to prove
    $H^{1}(C,K_{C}^{\frac{r-1}{2}}\otimes\mathcal{V}_{r-1}\otimes K_{C}^{\frac{1}{2}})\neq 0.$
    Since $K_{C}^{\frac{r-2}{2}}$ is a subbundle of $\mathcal{V}_{r-1}$, there is an exact sequence
    $0\to K_{C}^{\frac{r-1}{2}}\to\mathcal{V}_{r-1}\otimes K_{C}^{\frac{1}{2}}\to\mathcal{Q}_{1}\to 0,$
    where $\mathcal{Q}_{1}$ is a subbundle of rank $r-2$, and $\mathcal{F}_{r-1}^{\bullet}/K_{C}^{\frac{r-1}{2}}$ gives a filtration on $\mathcal{Q}_{1}$, which is Griffiths transverse with respect to the quotient connection. Since an exact sequence of locally free sheaves is again exact after tensoring a locally free sheaf, there is also an exact sequence
    $0\to K_{C}^{r-1}\to K_{C}^{\frac{r-1}{2}}\otimes\mathcal{V}_{r-1}\otimes K_{C}^{\frac{1}{2}}\to K_{C}^{\frac{r-1}{2}}\otimes\mathcal{Q}_{1}\to 0,$
    which gives a long exact sequence
    $\dots\to H^{1}(C,K_{C}^{r-1})\to H^{1}(C,K_{C}^{\frac{r-1}{2}}\otimes\mathcal{V}_{r-1}\otimes K_{C}^{\frac{1}{2}})\to H^{1}(C,K_{C}^{\frac{r-1}{2}}\otimes\mathcal{Q}_{1})\to 0.$
    Since $H^{1}(C,K_{C}^{r-1})=0$ for $r\geq 3$, we have
    $H^{1}(C,K_{C}^{\frac{r}{2}}\otimes\mathcal{V}_{r-1}\otimes K_{C}^{\frac{1}{2}})\cong H^{1}(C,K_{C}^{\frac{r-1}{2}}\otimes\mathcal{Q}_{1}).$
    If $r=3$, then $\mathcal{Q}_{1}=\mathcal{O}_{C}$, and thus
    $H^{1}(C,K_{C}^{\frac{r}{2}}\otimes\mathcal{V}_{r-1}\otimes K_{C}^{\frac{1}{2}})\cong H^{1}(C,K_{C})\cong k.$
    If $r>3$, then $K_{C}^{\frac{r-3}{2}}$ is a subbundle of $\mathcal{Q}_{1}$. If we use $\mathcal{Q}_{2}$ to denote the quotient of $\mathcal{Q}_{1}$ by $K_{C}^{\frac{r-3}{2}}$, then there is an exact sequence
    $0\to K_{C}^{r-2}\to K_{C}^{\frac{r-1}{2}}\otimes\mathcal{Q}_{1}\to K_{C}^{\frac{r-1}{2}}\otimes\mathcal{Q}_{2}\to 0$
    Since $r$ is odd, automatically we have $r>4$, and the long exact sequence gives the following isomorphism
    $H^{1}(C,K_{C}^{\frac{r-1}{2}}\otimes\mathcal{V}_{r-1}\otimes K_{C}^{\frac{1}{2}})\cong H^{1}(C,K_{C}^{\frac{r-1}{2}}\otimes\mathcal{Q}_{2}).$
    By repeating above process finite times, we can prove that
    $H^{1}(C,K_{C}^{\frac{r-1}{2}}\otimes\mathcal{V}_{r-1}\otimes K_{C}^{\frac{1}{2}})$ $\cong H^{1}(C,K_{C}^{\frac{r-1}{2}}\otimes\mathcal{Q}_{r-2})$
    $\cong H^{1}(C,K_{C})$
    $\cong k.$
  2. $(2)$
    If $r$ is even, then by induction hypothesis there exists an oper $(\mathcal{V}_{r-1},\nabla_{r-1})$ such that $\mathcal{V}_{r-1}$ is indecomposable and the underlying vector bundle of the associated Higgs bundle is
    $K_{C}^{-\frac{(r-2)}{2}}\oplus\dots\oplus\mathcal{O}_{C}\oplus\dots\oplus K_{C}^{\frac{r-2}{2}}.$
    By the same argument, it suffices to prove
    $H^{1}(C,K_{C}^{\frac{r-1}{2}}\otimes\mathcal{V}_{r-1}\otimes K_{C}^{\frac{1}{2}})\neq 0,$
    which can also be done by the same argument.
4.3.2. The construction of tensor product type counter-example
Let $(\mathcal{V}_{\operatorname{unif}},\nabla_{\operatorname{unif}})$ be uniformizing flat bundle constructed in Lemma 4.3.1. Then $0\subset K_{C}^{\frac{1}{2}}\subset\mathcal{V}$ is the unique Simpson filtration on $(\mathcal{V}_{\operatorname{unif}},\nabla_{\operatorname{unif}})$ such that the associated graded Higgs bundle is the uniformizing Higgs bundle.
Since $(\mathcal{V}_{\operatorname{unif}},\nabla_{\operatorname{unif}})$ is of rank two, we have the $p$-curvature of $\nabla_{\operatorname{unif}}$ is nilpotent of exponent $\leq 1$, and thus $(\mathcal{E}_{2},\theta_{2})=C(\mathcal{V}_{\operatorname{unif}},\nabla_{\operatorname{unif}})$ is a well-defined stable Higgs bundle, as $(\mathcal{V}_{\operatorname{unif}},\nabla_{\operatorname{unif}})$ is $\nabla$-stable.
Let $\mathcal{E}_{1}$ be the Frobenius pushforward of $K_{C}^{\frac{1-p}{2}}$. Then by Theorem 4.1.2 and Corollary 4.1.1 we have $\mathcal{E}_{1}$ is a stable rank $p$ bundle of degree zero. Moreover, by Lemma 4.1.2 the unique Simpson filtration on $F^{*}\mathcal{E}_{1}$ is given by its canonical filtration, and the associated graded Higgs bundle is
$(K_{C}^{\otimes\frac{1-p}{2}}\oplus\dots\oplus\mathcal{O}_{C}\oplus\dots\oplus K_{C}^{\otimes\frac{p-1}{2}},\bigoplus_{i=1}^{p-1}\theta_{i})$
where $\theta_{i}\colon K_{C}^{\otimes i-1}\to K_{C}^{\otimes i}$ is an isomorphism for each $i=\frac{1-p}{2},\dots,\frac{p-1}{2}$.
The Higgs bundle $(\mathcal{E}_{1}\otimes\mathcal{E}_{2},\operatorname{id}\otimes\theta_{2})$ is stable since $(F^{*}F_{*}K_{C}^{\frac{1-p}{2}},\nabla_{\operatorname{can}})\otimes(\mathcal{V}_{\operatorname{unif}},\nabla_{\operatorname{unif}})$ is $\nabla$-stable, and it initiates the following Higgs-de flow
[…]
where $(\mathcal{E}_{p},\theta_{p})=(K_{C}^{\otimes\frac{1-p}{2}}\oplus\dots\oplus\mathcal{O}_{C}\oplus\dots\oplus K_{C}^{\otimes\frac{p-1}{2}},\bigoplus_{i=1}^{p-1}\theta_{i})\otimes(K_{C}^{\frac{1}{2}}\oplus K_{C}^{-\frac{1}{2}},\theta)$ and $\mathcal{N}^{\bullet}$ is the filtration given by the tensor product of canonical filtration on $F^{*}F_{*}K_{C}^{\frac{1-p}{2}}$ and Simpson filtration on uniformizing flat bundle.
This gives a tensor product type counter-example of Lan-Sheng-Zuo conjecture , since by Lemma 4.3.2 we have $(\mathcal{E}_{p},\theta_{p})$ is nilpotent of exponent $=p$, and we cannot find any other Simpson filtration on $(F^{*}F_{*}K_{C}^{\frac{1-p}{2}},\nabla_{\operatorname{can}})\otimes(\mathcal{V}_{\operatorname{unif}},\nabla_{\operatorname{unif}})$ such that the associated graded Higgs bundle has smaller exponent of nilpotency, since $(\mathcal{E}_{p},\theta_{p})$ is polystable.
Lemma 4.3.2.
For any $k\in\operatorname{\mathbb{Z}}_{>0}$,
$(\mathcal{E}_{k},\theta_{k})=(K_{C}^{\otimes\frac{1-k}{2}}\oplus\dots\oplus\mathcal{O}_{C}\oplus\dots\oplus K_{C}^{\otimes\frac{k-1}{2}},\bigoplus_{i=1}^{k-1}\theta_{i})\otimes(K_{C}^{\frac{1}{2}}\oplus K_{C}^{-\frac{1}{2}},\theta)$
is a polystable Higgs bundle which is nilpotent of exponent $=k$.
Proof.
In order to prove $(\mathcal{E}_{k},\theta_{k})$ is a polystable Higgs bundle with nilpotent of exponent $=k$, we claim that $(\mathcal{E}_{k},\theta_{k})=(\mathcal{E}_{k}^{\prime},\theta_{k}^{\prime})\oplus(\mathcal{E}^{\prime\prime}_{k},\theta_{k}^{\prime\prime})$, where $\mathcal{E}^{\prime}_{k}$ is a stable rank $k+1$ Higgs bundle of degree zero with nilpotent of exponent $=k$, and $\mathcal{E}^{\prime\prime}_{k}$ is a stable rank $k-1$ Higgs bundle of degree zero with nilpotent of exponent $=k-2$.
The case $k=1$ is trivial. For the case $k=2$, we may write the Higgs field $\theta_{2}$ on $\mathcal{E}_{2}$ in terms of matrix as follows:
$\begin{pmatrix}0&1&1&0\\ 0&0&0&1\\ 0&0&0&1\\ 0&0&0&0\\ \end{pmatrix}\sim\begin{pmatrix}0&0&1&0\\ 0&0&0&1\\ 0&0&0&0\\ 0&0&0&0\\ \end{pmatrix}.$
Thus
$(\mathcal{E}_{2},\theta_{2})=(K_{C}\oplus\mathcal{O}_{C}\oplus\mathcal{T}_{C},\theta\oplus\theta^{\vee})\oplus(\mathcal{O}_{C},0),$
which shows the claim holds for $k=2$.
Suppose the claim holds for the $(\mathcal{E}_{k-1},\theta_{k-1})$ case and write
$(\mathcal{E}_{k-1},\theta_{k-1})=(\mathcal{E}_{k-1}^{\prime},\theta_{k-1}^{\prime})\oplus(\mathcal{E}_{k-1}^{\prime\prime},\theta_{k-1}^{\prime\prime}).$
Note that there is an exact sequence of Higgs bundles
$0\to(K_{C}^{-\frac{1}{2}}\oplus K_{C}^{\frac{1}{2}},\theta)\otimes(K_{C}^{\frac{k-1}{2}},0)\to(\mathcal{E}_{k},\theta_{k})\to(\mathcal{E}_{k-1},\theta_{k-1})\otimes(K_{C}^{-\frac{1}{2}},0)\to 0$
Moreover, it cannot split since $(\mathcal{E}_{k-1},\theta_{k-1})\otimes(K_{C}^{-\frac{1}{2}},0)$ is not $\theta_{k}$-invariant, which is violated by isomorphisms
$\mathcal{E}_{k-1}\otimes K_{C}^{-\frac{1}{2}}\supset K_{C}^{\frac{k-4}{2}}$ $\to K_{C}^{\frac{k-2}{2}}\subset(K_{C}^{-\frac{1}{2}}\oplus K_{C}^{\frac{1}{2}})\otimes K_{C}^{\frac{k-1}{2}}$
$\mathcal{E}_{k-1}\otimes K_{C}^{-\frac{1}{2}}\supset K_{C}^{\frac{k-2}{2}}$ $\to K_{C}^{\frac{k}{2}}\subset(K_{C}^{-\frac{1}{2}}\oplus K_{C}^{\frac{1}{2}})\otimes K_{C}^{\frac{k-1}{2}}.$
Since the components of Higgs field are given by isomorphisms, we can choose an appropriate basis of $K_{C}^{\frac{k-2}{2}}\oplus K_{C}^{\frac{k-2}{2}}\subset(\mathcal{E}_{k},\theta_{k})$ such that
$(\mathcal{E}_{k},\theta_{k})\cong\left((\mathcal{E}_{k-1}^{\prime}\otimes K_{C}^{-\frac{1}{2}})\oplus K_{C}^{\frac{k}{2}},\theta^{\prime}_{k}\right)\oplus\left((\mathcal{E}_{k-1}^{\prime\prime}\otimes K_{C}^{-\frac{1}{2}})\oplus K_{C}^{\frac{k-2}{2}},\theta^{\prime\prime}_{k}\right).$
Moreover, both $(\mathcal{E}_{k-1}^{\prime}\otimes K_{C}^{\frac{1}{2}})\oplus K_{C}^{\frac{k}{2}}$ and $(\mathcal{E}_{k-1}^{\prime\prime}\otimes K_{C}^{-\frac{1}{2}})\oplus K_{C}^{\frac{k-2}{2}}$ are stable of degree zero and the Higgs fields are of desired nilpotency. This completes the proof. ∎
5. Lan-Sheng-Zuo conjecture on higher dimensional varieties
In this section we prove a case of the Lan-Sheng-Zuo conjecture on higher dimensional varieties.
Proposition 5.1.
Let $X$ be a smooth $n$-dimensional projective variety over $k$ and $\omega$ be an ample polarization on $X$ such that $\mu_{\omega,\max}(\Omega_{X}^{1})\leq 0$. Then the Lan-Sheng-Zuo conjecture holds on $X$.
Proof.
Let $(\mathcal{E},\theta)$ be a semistable Higgs bundle of exponent $\leq p-1$ and $(\mathcal{V},\nabla)=C^{-1}(\mathcal{E},\theta)$. Then $(\mathcal{V},\nabla)$ is a $\nabla$-semistable flat bundle. Now it suffices to prove that $\mathcal{V}$ is semistable as a vector bundle. If so, then by Theorem 4.1.1 we have $\mathcal{V}$ is strongly semistable, and thus we can always choose trivial filtration to run the Higgs-de Rham flow.
Suppose $\mathcal{V}$ is unstable and its Harder-Narasimhan filtration is given by
$0=\mathcal{V}_{0}\subset\mathcal{V}_{1}\subset\dots\subset\mathcal{V}_{s-1}\subset\mathcal{V}_{s}=\mathcal{V},$
where $\mathcal{G}_{i}=\mathcal{V}_{i}/\mathcal{V}_{i-1}$ is $\omega$-semistable bundle for each $i$ with strictly decreasing slopes. Since $(\mathcal{V},\nabla)$ is $\nabla$-semistable, we must have the maximally destabilizer $\mathcal{V}_{1}$ is not $\nabla$-invariant. Choose $\ell\in\operatorname{\mathbb{Z}}_{>1}$ such that $\nabla\left(\mathcal{V}_{1}\right)\subseteq\mathcal{V}_{\ell}$ but $\nabla\left(\mathcal{V}_{1}\right)\nsubseteq\mathcal{V}_{\ell-1}$. Then $\nabla$ induces a non-zero morphism $\mathcal{G}_{1}\to\mathcal{G}_{\ell}\otimes\Omega_{X}^{1}$.
Suppose the maximal destabilizer of $\Omega_{X}^{1}$ is $(\Omega_{X}^{1})_{1}$. Then by Theorem 4.1.1 we have both $(\Omega_{X}^{1})_{1}$ and $\mathcal{G}_{i}$ are strongly semistable, and thus the tensor product $\mathcal{G}_{\ell}\otimes(\Omega_{X}^{1})_{1}$ is semistable. As a consequence, we have
$\mu_{\omega,\max}(\mathcal{G}_{\ell}\otimes\Omega_{X}^{1})$ $\leq\mu_{\omega}(\mathcal{G}_{\ell})+\mu_{\omega,\max}(\Omega_{X}^{1})$
$\leq\mu_{\omega}(\mathcal{G}_{\ell})$
$<\mu_{\omega}(\mathcal{G}_{1}).$
This shows there is no non-zero morphism between $\mathcal{G}_{1}$ and $\mathcal{G}_{\ell}\otimes\Omega_{X}^{1}$ for $\ell\in\operatorname{\mathbb{Z}}_{>1}$, a contradiction. ∎